OpenStudy (anonymous):

Help Please! Given the following perfect square trinomial, fill in the missing term. x2 − 16x + ____

OpenStudy (anonymous):

what is half of 16?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@satellite73 8 I never learned this so I need help learning how to

OpenStudy (anonymous):

half of 16 is 8 and \(8^2=64\) that is your answer

OpenStudy (anonymous):

\[x^2-16+64=(x-8)^2\] a perfect square

OpenStudy (anonymous):

wanna try another one?

OpenStudy (anonymous):

sure @satellite73